1. An aqueous solution of calcium hydroxide is standardized by titration with a 0.200 M solution of perchloric acid.

If 19.8 mL of base are required to neutralize 13.0 mL of the acid, what is the molarity of the calcium hydroxide solution?

__M calcium hydroxide

2. An aqueous solution of calcium hydroxide is standardized by titration with a 0.143 M solution of nitric acid.

If 13.0 mL of base are required to neutralize 10.1 mL of the acid, what is the molarity of the calcium hydroxide solution?

___M calcium hydroxide
1
Expert's answer
2019-10-18T08:06:00-0400

Solution.

1.

Write down the reaction equation. 

"Ca(OH)2 + 2HClO4 = Ca(ClO4)2 + 2H2O"

Since we need to find the molarity of calcium hydroxide, we can leave the molarity of chloric acid. Translate ml to L.

"C(HClO4)*V(HClO4) = C(Ca(OH)2)*V((CaOH)2)"

"C(Ca(OH)2) = \frac{C(HClO4)*V(HClO4)}{V(Ca(OH)2)}"

C(Ca(OH)2) = 0.131 M

2.

Write down the reaction equation. 

"Ca(OH)2 + 2HNO3 = Ca(NO3)2 + 2H2O"

Since we need to find the molarity of calcium hydroxide, we can leave the molarity of chloric acid. Translate ml to L.

"C(HNO3)*V(HNO3) = C(Ca(OH)2)*V((CaOH)2)"

"C(Ca(OH)2) = \frac{C(HNO3)*V(HNO3)}{V(Ca(OH)2)}"

C(Ca(OH)2) = 0.1111 M

Answer:

1.

C(Ca(OH)2) = 0.131 M

2.

C(Ca(OH)2) = 0.111 M